The magnitude of an earthquake is given by ๐‘€ = log ๐ผ


where M is the Richter scale value of an earthquake, and I is the intensity of the earthquake. An earthquake of magnitude 6.4 is measured and the following aftershock has a magnitude or 4.1.


Determine algebraically how many times more intense, to the nearest whole, the earthquake is compared to the aftershock.


1
Expert's answer
2021-11-22T13:52:22-0500
"M=\log_{10}I=>I=10^M"

"\dfrac{I_E}{I_{after}}=\dfrac{10^{M_E}}{10^{M_{after}}}=10^{M_E-M_{after}}"

"=10^{6.4-4.1}=10^{2.3}\approx200"

An earthquake is 200 times more intense in comparison with the following aftershock.